OpenStudy (anonymous):

solve the difference equation yn+1=-.8Yn+90

OpenStudy (anonymous):

got any answers to this

OpenStudy (anonymous):

yn=111.25 if wrong correct me

OpenStudy (anonymous):

forget yn+1

OpenStudy (anonymous):

just 1=-0.8yn+90

OpenStudy (anonymous):

do opposite of add its subtract

OpenStudy (anonymous):

so 1-90=-89

OpenStudy (anonymous):

-89/-0.8=111.25

ganeshie8 (ganeshie8):

\[\begin{align}y_{n+1} &= -0.8Y_n +90\\~\\&=-0.8\left(-0.8Y_{n-1} + 90\right) + 90 = (-0.8)^2Y_{n-1} + 90(1 -0.8)\\~\\&=(-0.8)^2(-0.8Y_{n-2}+90) + 90(1 -0.8) = (-0.8)^3Y_{n-2} + 90(1 -0.8 + (-0.8)^2)\\~\\~\\&=\end{align}\]

ganeshie8 (ganeshie8):

see any pattern ?

OpenStudy (anonymous):

omg, i figured it out! thank you!

ganeshie8 (ganeshie8):

contunuing further you would get : \[\large Y_{n+1} = (-0.8)^{n}Y_1 + 90\left(\dfrac{1-(-0.8)^{n}}{1-(-0.8)}\right) \]

OpenStudy (anonymous):

right